OpenStudy (anonymous):

Write in simplest notation and evaluate: (16^2)^3/(2^5)^4? (i keep getting 512 but the answer says 16, can anyone see where i went wrong?)

OpenStudy (anonymous):

i dont know where you went wrong but you have to make sure you follow the order of operations. so square 16 first and then raise that number to the 3rd power (you get 16777216)... then raise 2 to the 5th power (you get 32) and then raise that to the 4th power (you get 1048576). then divide 16777216 by 1048576 and you get 16

OpenStudy (anonymous):

numerator is \[16^6=2^{24}\] denominator is \[2^{20}\] giving \[\frac{2^{24}}{2^{20}}=2^4=16\]

OpenStudy (anonymous):

or i guess you could raise all this stuff to those crazy big numbers. i wouldn't for sure

OpenStudy (anonymous):

your way works as well.. i just find my way easier if using a calculator

OpenStudy (anonymous):

btw i hope it is clear that since \[16=2^4\] then \[16^6=2^{4\times 6}=2^{24}\]